OpenStudy (anonymous):

Help! Will give medal! The steps below show the incomplete solution to find the value of n for the equation 4n - 2n + 4 = -1 + 17 Step 1: 4n - 2n + 4 = 16 Step 2: 4n - 2n + 4 = 16 Step 3: 2n + 4 = 16 Which of these is most likely the next step A) 2n = 12 B) 2n = 4 C) 2n = 8 D) 2n = 20 I believe it is C but can someone double check?

OpenStudy (anonymous):

You want 2n to be by itself on one side. to do that you need to move 4 to the other side. subtract 4 on both sides. 2n=12

OpenStudy (kendricklamar2014):

Like @GretaKnows said:\[2n + 4-4 = 16 - 4\] Which equals: \[2n = 12 ~(Choice ~ A)\]